### v3

# Beyond the Kolmogorov Barrier: A Learnable Weighted Hybrid Autoencoder for Model Order Reduction

Abstract
Representation learning for high-dimensional, complex physical systems aims to identify a low-dimensional intrinsic latent space, which is crucial for reduced-order modeling and modal analysis. To overcome the well-known Kolmogorov barrier, deep autoencoders (AEs) have been introduced in recent years, but they often suffer from poor convergence behavior as the rank of the latent space increases. To address this issue, we propose the learnable weighted hybrid autoencoder, a hybrid approach that combines the strengths of singular value decomposition (SVD) with deep autoencoders through a learnable weighted framework. We find that the introduction of learnable weighting parameters is essential — without them, the resulting model would either collapse into a standard POD or fail to exhibit the desired convergence behavior. Interestingly, we empirically find that our trained model has a sharpness thousands of times smaller compared to other models. Our experiments on classical chaotic PDE systems, including the 1D Kuramoto-Sivashinsky and forced isotropic turbulence datasets, demonstrate that our approach significantly improves generalization performance compared to several competing methods. Additionally, when combining with time series modeling techniques (e.g., Koopman operator, LSTM), the proposed technique offers significant improvements for surrogate modeling of high-dimensional multi-scale PDE systems.
keywords:
Reduced-order modeling , Dimensionality Reduction , POD , Deep Learning, Surrogate Modeling
[…]
[1]organization=Department of Mechanical, Aerospace, and Nuclear Engineering, Rensselaer Polytechnic Institute, addressline=110 8th St, city=Troy, postcode=12180, state=New York, country=USA
[…]
[2]organization=Scientific Computation Research Center, Rensselaer Polytechnic Institute, addressline=110 8th St, city=Troy, postcode=12180, state=New York, country=USA
1 Introduction
Computational fluid dynamics involves solving large dynamical systems with millions of degrees of freedom, resulting in significant computational overhead. In order to alleviate this problem, reduced-order modeling [1] is widely used, which uses a smaller number of modes to provide an approximate solution at a lower computational expense. A crucial step in reduced-order modeling is the projection of system states to a reduced latent space. The projection error on the latent space represents the overhead error that affects the solution accuracy throughout the time-stepping scheme. Linear dimensionality reduction techniques such as Proper Orthogonal Decomposition (POD) [2, 3, 4] are often used to create efficient representations of large-scale systems by projecting the solution manifold onto the space spanned by a set of linear orthonormal basis. Advances in deep learning techniques, such as deep autoencoders (AE) [5], capture intrinsic non-linear features for better compression and retrieval of high-fidelity information, outperforming POD in lower ranks, thus overcoming the well-known Kolmogorov barrier [6, 7]. Milano and Koumoutsakos [8] highlighted one of the first works to utilize a fully connected autoencoder to reconstruct the flow field, offering better performance compared to POD. Further studies have reported the usage of convolutional neural networks on 2D or 3D flow fields [9, 10, 11]. Such methods have been adopted in the fluid dynamics community to obtain a nonlinear model order reduction [12], but they do not provide rapid error convergence at higher rank. Recently, there have also been studies on using the hybrid approach [13, 14] combining POD with deep learning, by passing the latent space produced by POD to a neural network to find the corrections required to enhance reconstruction. These hybrid techniques have proven to enhance reconstruction beyond the capabilities of vanilla autoencoders. Following a similar path, here we propose a novel dimensionality reduction technique that combines traditional dimensionality reduction technique, i.e., POD, with deep learning techniques in a weighted manner at the encoder and decoder stage, where such weights of hybridization are also learned from data, to achieve a more effective and flexible dimensionality reduction. We also compare this approach with a straightforward hybrid technique, where a direct sum of POD and AE is utilized to construct the autoencoder, demonstrating the need of using learnable weighting parameters between POD and AE. Interestingly, our proposed approach obtains flat minima as opposed to other approaches, which contributes to the improved generalization and noise robustness.
Building on our proposed framework, we also demonstrate the application of our framework to surrogate modeling tasks, particularly in the context of system evolution. We explore two distinct applications of our framework. First, we integrate our framework with the Koopman operator, where our framework facilitates the mapping of Koopman embedding back to physical space, while learning a linear forward model to evolve the system states in a reduced space. Second, we leverage long-short-term memory (LSTM) for system evolution [15]. LSTMs are usually trained on reduced states, obtained via POD [16, 17] or a non-linear autoencoder [18, 19] to serve as a surrogate for the temporal evolution of latent states. The computational costs can be prohibitive when using LSTMs on full order states, which can typically be in the range of millions for complex scenarios in fluid dynamics. By integrating our improved dimensionality reduction technique with a forward model for the system (linear or non-linear), we show that the enhanced latent space leads to more accurate prediction of the system dynamics and provide insights into the individual error contributions from the dimensionality reduction and time series model. This underscores the fact that the quality of the reduced representation is the primary bottleneck in achieving highly accurate predictions, as suboptimal latent spaces inherently limit the effectiveness of any downstream models.
The remainder of the manuscript is structured as follows. In Section 2 we present the details of our hybrid dimensionality reduction technique, Koopman operator used for long term forecast and nonlinear surrogate for learning PDE dynamics. Section 3 describes the datasets used in each of the tasks, network architecture, and training hyper-parameters. Section 4 provides a comprehensive evaluation of the proposed framework with respect to pure dimensionality reduction tasks and other downstream applications, comparing its performance against other techniques, and examining some key properties related to sharpness and noise robustness. Finally, Section 5 concludes the paper.
2 Methodology
2.1 Dimensionality Reduction
Without loss of generality, we begin by sampling a general vector-valued spatial-temporal field $u(x,t)\in\mathbb{R}^{Q}$ on a fixed mesh with $N$ cells, where $(x,t)$ is the space-time coordinate. At each time $t$, a cell-centered snapshot sample is a matrix $\mathbf{x}\in\mathbb{R}^{N\times Q}$. In the current framework, we start with two encoders:
  1. 1.
    POD based encoder using $r$-dominant left singular vectors from the matrix consisting of stacked flattened columns of $\mathbf{x}$, denoted as $\phi_{\textrm{POD}}$.
  2. 2.
    the neural network encoder with output dimension as $r$, denoted as $\phi_{\textrm{NN}}$. As shown in eq. 1, the latent state $\mathbf{z}\in\mathbb{R}^{r}$ is obtained by a weighted sum of POD projection and the output of encoder,
    $\mathbf{z}=(\mathbf{1}-\mathbf{a})\phi_{\textrm{POD}}(\mathbf{x})+\mathbf{a}\phi_{\textrm{NN}}(\mathbf{x}),$ (1)
    where $\mathbf{1}\in\mathbb{R}^{r}$ is a vector of ones, and $\mathbf{a}\in\mathbb{R}^{r}$ is a vector of learnable weights.
Next for the decoder part, we project the latent state $\mathbf{z}$ back to the reconstructed system state following eq. 2,
$\hat{\mathbf{x}}=\psi_{\textrm{POD}}(\mathbf{z})(\mathbf{I}-\mathbf{B})+\psi_{\textrm{NN}}(\mathbf{z})\mathbf{B},$ (2)
where $\mathbf{I}$ is $Q\times Q$ identity matrix, $\mathbf{B}=\textrm{diag}(\mathbf{b})$, $\mathbf{b}\in\mathbb{R}^{Q}$, $\psi_{\textrm{POD}}$, and $\psi_{\textrm{NN}}$ are POD decoder and NN decoder, respectively. In addition to the parameters of NN encoder and decoder, both $\mathbf{a}\in\mathbb{R}^{r}$ and $\mathbf{b}\in\mathbb{R}^{Q}$ are trainable through gradient-based optimization as well. Hence, we name the above framework as learnable weighted hybrid autoencoder. It is important to note that such NN can be either fully-connected or convolutional.
[…]
To emphasize the role of learnable weight $\mathbf{a}$ and $\mathbf{B}$, we also implement a straightforward hybrid approach [22], which simply adds the latent states from POD and NN. Similarly, the state of the reconstructed system is the sum of the output of the POD decoder and the NN decoder. We name this hybrid approach as simple hybrid autoencoder. The difference of two approaches is further illustrated in fig. 1. As we shall see in the following sections, this comparison underscores the importance of using a weighted blend of the encoded and decoded spaces, as without it, the improvement over POD is at most incremental.
Figure 1: Architecture of the simple hybrid autoencoder and learnable weighted hybrid autoencoder.
2.2 Koopman Forecasting
Koopman theory [23] states that any nonlinear dynamical system can be linearized by lifting into a (possible infinite-dimensional) space of observable functions. Hence, given a continuous dynamical system $\dot{\mathbf{x}}=\mathbf{f}(\mathbf{x})$, for any observable $\xi$, one defines the Koopman operator $\mathcal{K}$ as:
$\xi(\mathbf{x}(t+\Delta t))=\mathcal{K}\xi(\mathbf{x}(t)),$ (3)
where $t$ refers to the time and $\Delta t$ is the time step size. Although the idea of linear evolution is attractive and desirable in most scenarios, the effectiveness of Koopman theory is often limited by the subspace chosen.
To overcome these challenges, Lange et al. [24] proposed Koopman forecasting,
$\xi(\mathbf{x}(t))=\begin{bmatrix}\cos(\vec{\mathbf{\omega}}t)\\ \sin(\vec{\mathbf{\omega}}t)\end{bmatrix}:=\Omega(\vec{\mathbf{\omega}}t),$ (4)
where $\vec{\omega}$ refers to the frequency vector. With these assumptions the system state at any time $t$ is given by
$\mathbf{x}(t)=\psi_{\theta}(\Omega(\vec{\omega}t)).$ (5)
We solve the following minimization problem to learn the non-linear mapping $\psi_{\theta}$ and frequencies $\vec{\omega}$, where $\psi_{\theta}$ is parametrized by $\theta$ which includes the neural network’s weights, biases and the learnable weighting parameters :
$\min_{\vec{\omega},\theta}\sum_{i=1}^{T}\left\|\mathbf{x}_{i}-\psi_{\theta}(\Omega(\vec{\omega}i\Delta t))\right\|^{2},$ (6)
where $i$ indicates the timestep index and $\Delta t$ is the timestep size between consecutive snapshots. We utilize our proposed hybrid framework to learn the mapping from observable space to the physical state, while keeping their time evolution strategy unchanged, taking the same form as in eq. 2
$\psi_{\theta}(\Omega(\vec{\omega}t))=\psi_{\textrm{POD}}(\Omega(\vec{\omega}t))(\mathbf{I}-\mathbf{B})+\psi_{\textrm{NN}}(\Omega(\vec{\omega}t))\mathbf{B},$ (7)
with $\mathbf{B}$ initialized to zeros during training. It should be noted that this problem involves only the learning of a decoder jointly with the Koopman model for time evolution.
2.3 Surrogate Modeling for Time-Dependent PDEs
Our surrogate modeling strategy involves the usage of the aforementioned techniques to obtain a latent representation and LSTM to determine its evolution over time. Training is divided into two stages. The dimensionality reduction framework is first trained to obtain a low-dimensional embedding of the system states, which subsequently serves as the training data for the time series prediction model to evolve the system in time. The advantage of training them separately is that it allows us to clearly demarcate the contribution of each stage to the overall surrogate model, which is particularly important as the crux of the current study is the introduction of a novel framework for dimensionality reduction.
LSTM networks are a type of recurrent neural network (RNN) designed to address the vanishing gradient problem in standard RNNs. LSTMs incorporate a gating mechanism to selectively retain and propgate information over long sequences, making them suitable for modeling sequential and time-dependent data. They are auto-regressive in nature, using data from the past $k$ timesteps to predict the next state. To predict $x_{i+1}$, the model uses information from $x_{i},x_{i-1},\ldots,x_{i-k+1}$, effectively capturing temporal dependencies in multiple time steps. $k$ is referred to as the look-back window and is usually tuned based on the problem and the dataset at hand. In the current study, we set the look-back window to a value of 10 for every dataset. This value was found to be optimal in our analysis. Other architectural details of the LSTM will be discussed on a case-by-case basis.
3 Datasets and model setup
3.1 Chaotic System
3.1.1 Kuramoto-Sivashinsky (KS)
The governing equation of the KS equation is $\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}+\frac{\partial^{2}u}{\partial x^{2}}+\frac{\partial^{4}u}{\partial x^{4}}=0$, where $x\in[0,L_{x})$, $L_{x}=64\pi$, and periodic boundary conditions are assumed. The initial conditions comprise the sum of ten random sine and cosine waves $u(x,0)=\sum_{k=1}^{10}A_{k}\left(\sin\left(\frac{2\pi n_{k}x}{L_{x}}+\phi_{k}\right)+\cos\left(\frac{2\pi n_{k}x}{L_{x}}+\phi_{k}\right)\right)$, where $\forall k\in\{1,\ldots,10\},A_{k}\sim U(-1,1),\phi_{k}\sim U(0,2\pi)$ and wavenumber $n_{k}\sim U\{1,\ldots,6\}$, respectively. We employ Fourier decomposition in space and 2nd order Crank-Nicolson Adam-Bashforth semi-implicit finite difference scheme for temporal discretization. The data is generated for around 2000 time steps with a timestep of 0.01. Grid sizes of 512, 1024 and 2048 is used for the simulation. Some initial snapshots of the simulation that correspond to the transient phase are ignored.
For the NN encoder, we use a feedforward neural network with one hidden layer as shown in fig. 2, $2r$ number of neurons in the hidden layer. Hyperbolic tangent function is used for activation in all layers except the last linear layer. The NN decoder is symmetric to the encoder. After shuffling, the snapshot data is split into training and testing with a 7:3 ratio. Before training, we standardize the data by subtracting the mean and dividing by the standard deviation. All models are trained for 40k epochs with a batch size of 64.
Figure 2: NN-based encoder-decoder utilizing fully connected neural networks.
Figure 3: NN-based encoder-decoder utilizing convolutional neural networks.
3.1.2 Homogeneous isotropic turbulence (HIT)
Direct numerical simulation data of forced homogeneous isotropic turbulence is obtained from the Johns Hopkins turbulence database [25]. The data set is generated by solving the forced Navier-Stokes equation on a periodic cubic box using the pseudospectral method. We interpolate the velocity field, $u(x,t)=(u_{x},u_{y},u_{z})\in\mathbb{R}^{3}$, from the original dataset of resolution $1024^{3}$ to grid sizes of $16^{3}$, $32^{3}$ and $64^{3}$ and extract the data from timestep 1 to 2048 with a stride of 16, resulting around 128 snapshots.
We choose a deep convolutional autoencoder [26] as the NN part of the proposed framework, as shown in fig. 3. The details of the architecture pertaining to the NN based encoder and decoder and the training hyper parameters are shown in table 1. Without shuffling, we systematically sample the dataset by selecting every alternate snapshot to curate the training and testing data, each comprising 64 snapshots.
Table 1: Architecture details and training parameters for 3D HIT dataset
Neural Network Architecture
Encoder Architecture Decoder Architecture
Component Details Component Details
Hidden Layers 4 Convolutional Layers Hidden Layers 4 Transpose Convolutional Layers
Filters 256, 512, 1024, 2048 Filters 2048, 1024, 512, 256
Bottleneck Layer Fully Connected Linear Layer Output Layer Fully Connected Linear Layer
Activation Function Swish (except bottleneck layer) Activation Function Swish (except output layer)
Dropout 0.4 (All layers except bottleneck) Dropout 0.4 (All layers except output)
Training Hyperparameters
Parameter Value
Epochs 2000
Batch Size 20
Optimizer Adam
Learning Rate $1\times 10^{-4}$
3.2 Koopman Forecasting Datasets
3.2.1 Traveling Wave
We demonstrate our framework on a traveling wave problem with a spatial dimensionality of 256. The spatio-temporal evolution of the wave is given by the following:
$u(x,t)=\mathcal{N}\left(x\mid\mu=100(\sin(0.01t)+1)+28,\sigma^{2}=10\right),$ (8)
where $\mathcal{N}$ is the probability density function of Gaussian distribution. We generate a trajectory of 100,000 timesteps, of which the initial 50,000 snapshots are used for training the Koopman-Decoder model, and the remaining 50,000 timesteps are held out for testing. The details of the non-linear part of the decoder and training parameters are provided in table 2. The Koopman model is parametrized by a single frequency, essentially evolving the dynamics in a latent space of rank 2.
Table 2: Decoder Architecture and Training Hyperparameters for the traveling wave problem
Decoder Architecture
Component Details
Initial Layer Fully Connected Linear Layer
Hidden Layers 4 Transpose Convolutional Layers
Channels 32, 16, 8, 4
Activation Function SiLU (except after linear and final layer)
Training Hyperparameters
Parameter Value
Optimizer Adam
Learning Rate $3\times 10^{-4}$
Batch Size 1280
Epochs 1000
3.2.2 Flow Over Cylinder
We utilize the time evolution data of vorticity in a two-dimensional flow over a cylinder at $Re\approx 100$. The initial 50 temporal snapshots are used in training the Koopman-Decoder model, and the remaining 100 are kept out for testing. The details of the non-linear part of the decoder and training parameters are provided in table 3. The Koopman model is parametrized by two frequencies, essentially evolving the dynamics in a latent space of rank 2.
Table 3: Decoder Architecture and Training Hyperparameters for the flow over cylinder problem
Decoder Architecture
Component Details
Initial Layer Fully Connected Linear Layer
Hidden Layers 4 Transpose Convolutional Layers
Channels 256, 128, 64, 32
Activation Function SiLU (except after linear and final layer)
Training Hyperparameters
Parameter Value
Optimizer Adam
Learning Rate $3\times 10^{-4}$
Batch Size 8
Epochs 1000
3.3 Surrogate Modeling Datasets
3.3.1 1D Viscous Burgers’
One dimensional Viscous Burgers’ equation is represented by the following partial differential equation.
$\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}=\nu\frac{\partial^{2}u}{\partial x^{2}},$ (9)
$u(x,0)=u_{0},\quad x\in[0,L],\quad u(0,t)=u(L,t)=0.$
The initial condition used is given by $u(x,0)=\frac{x}{1+\sqrt{\frac{1}{t_{0}}}\exp\left(Re\frac{x^{2}}{4}\right)}$. This form of the initial condition allows for an analytical solution of the form $u(x,t)=\frac{\frac{x}{t+1}}{1+\sqrt{\frac{t+1}{t_{0}}}\exp\left(Re\frac{x^{2}}{4(t+4)}\right)}$, where $t_{0}=e^{Re/8}$ and $Re=\frac{1}{\nu}$. Data generation is carried out by computing the analytical solution on a grid of 128 points for 100 timesteps with a terminal time of 2s. Model training utilizes trajectories of 19 different $Re$ values ranging from 100 to 1900 in steps of 100, with 100 time steps in each. For testing, we hold out trajectories of 13 different $Re$ values ranging from 50 to 2450 in steps of 200.
A fully connected network using the same architecture as outlined in section 3.1.1 is used for the NN based encoder and decoder. The dimensionality reduction framework was trained with a learning rate of $31e-$ 3, batch size of 32 for 500 epochs using the Adam optimizer. To enhance convergence, we used a cyclic learning rate scheduler, which dynamically adjusts the learning rate during training. A latent space dimension of 2 is chosen for this problem.
Following the training and convergence of the dimensionality reduction framework, the encoder is utilized to obtain the reduced-order representations. The resulting embeddings, denoted as $\mathbf{z}\in\mathbb{R}^{r}$, are subsequently augmented with the Reynolds number ($Re$), thereby extending the latent space dimension from $r=2$ to $r^{\prime}=3$. This enriched representation, $\tilde{\mathbf{z}}=[\mathbf{z},Re]$, serves as the input to the LSTM model, where the objective is to approximate the discrete-time mapping $\tilde{\mathbf{z}}_{i+1}=\mathcal{F}_{\theta}(\tilde{\mathbf{z}}_{i})$, parameterized by $\theta$ (LSTM weights and biases of the NN). At this stage, the emphasis shifts from state compression to modeling the underlying temporal dynamics in the reduced-order manifold. The LSTM is then trained to learn the evolution of $\tilde{\mathbf{z}}$, facilitating the construction of a surrogate model capable of forecasting trajectory evolution given initial conditions and governing parameters.
The LSTM consists of 2 hidden layers with 40 neurons. The model is trained with a batch size of 32 for 400 epochs with a learning rate of $10^{-3}$ using Adam optimizer along with a cyclic learning rate scheduler. Data scaling is not performed at any stage of the training. The task here is to learn a parametric surrogate model capable of capturing the system state evolution via a reduced representation.
3.3.2 2D Shallow Water
We extend the surrogate modeling to two dimensions by using inviscid shallow water equations of the form
$\frac{\partial(\rho\eta)}{\partial t}+\frac{\partial(\rho u\eta)}{\partial x}+\frac{\partial(\rho v\eta)}{\partial y}$ $=0,$ (10)
$\frac{\partial(\rho u)}{\partial t}+\frac{\partial}{\partial x}\left(\rho u^{2}+\frac{1}{2}\rho g\eta^{2}\right)+\frac{\partial(\rho uv)}{\partial y}$ $=0,$ (11)
$\frac{\partial(\rho v)}{\partial t}+\frac{\partial(\rho uv)}{\partial x}+\frac{\partial}{\partial y}\left(\rho v^{2}+\frac{1}{2}\rho g\eta^{2}\right)$ $=0.$ (12)
with $\eta$ representing the fluid column height, ($u,v$) referring to the depth averaged horizontal and vertical velocity of the fluid, $\rho$ being density of the fluid and $g$ is the acceleration due to gravity. The fluid density is kept constant at a value of 1.0. The governing equations are solved in a square domain of unit dimension, discretized using 64 grid points in each direction, periodic boundary conditions are enforced and initial conditions being
$\rho\eta(x,y,t=0)$ $=\exp\left(-\frac{(x-x_{0})^{2}}{0.005}-\frac{(y-y_{0})^{2}}{0.005}\right),$ (13)
$\rho u(x,y,t=0)$ $=0,$ (14)
$\rho v(x,y,t=0)$ $=0.$ (15)
The initial conditions refer to a Gaussian pulse in the fluid column height, where the parameters $x_{0}$ and $y_{0}$ denote the initial location of the Gaussian pulse. We generate 90 trajectories for training, each representing a different initial location of pulse with a timestep of 0.001 s. The simulation is run for a final time of 0.5 s with the states being saved every 5 timesteps, leaving us 100 timesteps in each trajectory. Other 10 trajectories are held for testing purposes.
A deep convolutional autoencoder is used for the NN part of the framework with architecture as outlined in section 3.1.2, with the only difference being that convolution and deconvolution operations were performed on two-dimensional data. The dimensionality reduction framework was trained first with a learning rate of $3\times 10^{-4}$, batch size of 24 for 500 epochs using Adam optimizer along with cyclic learning rate scheduler. We set the latent space rank to $r=6$. The full order states are scaled to zero mean and unit variance. Similar to the 1D Viscous Burgers case, we augment the latent space variable before training the LSTM model. Since this is a nonparametric PDE, the latent space variable here is augmented with the initial location of the pulse $\bar{x},\bar{y}$. This takes the LSTM input dimension from 6 to 8 for every trajectory. The LSTM consists of 3 hidden layers with 50 neurons. The model is trained with a batch size of 24 for 400 epochs with a learning rate of $10^{-3}$ using Adam optimizer along with a cyclic learning rate scheduler. The task here is to learn a surrogate model in reduced space given the initial condition and location of the pulse.
3.3.3 3D Viscous Burgers’
A straightforward extension of the 1D Viscous Burgers’ into three dimensions gives us the three-dimensional Viscous Burgers’ equation of the form
$\frac{\partial\mathbf{u}}{\partial t}=-b\frac{1}{2}\nabla\cdot(\mathbf{u}\otimes\mathbf{u})+\nu\nabla\cdot\nabla\mathbf{u},$ (16)
where $\mathbf{u}=(u,v,w)$ is the velocity field, $b$ is the advection parameter and $\nu$ is the diffusion parameter. Simulation data for the 3D Viscous Burgers’ equation is obtained from APEBench dataset [27]. The advection and diffusion parameters are set to a value of -1.5 and 1.5 respectively emulating a decaying dynamics. We generate the simulation data for 50 different trajectories, each with a different initial condition on a grid of size $32^{3}$ containing 101 temporal snapshots in each trajectory. For this dataset, we use the first 50 timesteps for training and the rest for testing purposes.
We use a deep convolutional autoencoder for the NN part of the framework with the same architecture as outlined in section 3.1.2. The dimensionality reduction framework was trained with a learning rate of 1e-4, batch size of 20 for 3000 epochs using Adam optimizer. We choose a six-dimensional latent space to represent the system. The full orders states are scaled to zero mean and unit variance prior to training.
For this problem, we do not augment the latent space with any control parameters prior to training the LSTM model. The LSTM consists of 3 hidden layers with 128 neurons. The learning hyperparameters are similar to the 2D Shallow Water case. In this case we evaluate the model’s ability to extrapolate in time using the latent representation, which can be challenging given the dynamics is decaying.
4 Results and Discussions
4.1 Dimensionality Reduction Performance
To demonstrate the effectiveness of the proposed approach, we compare our model against three other methods. POD, vanilla deep autoencoder (AE), simple hybrid autoencoder, as autoencoders for the two chaotic PDE datasets in sections 3.1.1 and 3.1.2 with varying rank $r$ and resolution $N$. Sixteen ensembles of the model are trained by setting different random seeds. From fig. 4, both of the two hybrid approaches perform better than AE and POD. However, simple approach does not maintain its convergence with increasing rank in contrast to our proposed approach. In addition, our approach shows an order-of-magnitude improvement in generalization as compared to any of the other methods. The simple and our learnable weighted hybrid approach give nearly the same test error at low ranks, but their gap increases multiple times with increasing rank. Surprisingly, such substantial improvement merely requires a negligible additional trainable parameters (i.e. $r+Q\ll N$), as shown in table 5. For this dataset the convergence of our approach follows a SVD-like convergence, while the simple approach has a behavior similar to AE.
Figure 4: Generalization performance of four models on 1D Kuramoto-Shivaskinsky dataset with varying rank $r$ and resolution $N$. X axis denotes the rank of the latent space and Y axis denotes the Mean Testing $L^{2}$ error. For each grid, rank, and method, the model is trained across 16 independent runs with varying random seed values. The results from these runs are averaged to compute the mean testing $L^{2}$ error. The convergence rate for POD and Learnable weighted hybrid approach is indicated in the legend as an exponent to $r$.
For the more challenging 3D HIT dataset, fig. 5 shows that our proposed approach continues to outperform the other three methods in terms of generalization, especially when the resolution increases (e.g., $32^{3},64^{3}$ as opposed to $16^{3}$). It is important to note that the simple approach shows little improvement over POD at resolution of $32^{3}$ or $64^{3}$ while our approach excels. Again, this highlights the important role of learnable weights in our hybrid approach. Additionally for this dataset, we study the impact of activation function on the reconstruction performance by utilizing ReLU activation function for the NN part of the auto-encoder. It is worth noting that the performance of both the AE and simple hybrid approaches exhibit noticeable variations when the activation function is changed. In contrast, our approach shows minimal changes in performance, demonstrating the robustness of the proposed framework. A further comparison among four models is shown in fig. 6, which visualizes the improvement in performance of our learnable weighted approach, especially on the small-scale turbulence. It is interesting to note that both POD and simple hybrid approach predicts a overly negative $u_{y}$ near the right middle. In addition, as shown in table 6, introducing such learnable weights has almost no overhead on the total number of parameters.
Figure 5: Generalization performance of four models on 3D homogeneous isotropic turbulence dataset with varying rank $r$ and resolution $N$. The testing $L^{2}$ error obtained using ReLU activation function is indicated using dashed lines. Similar to K-S case, 16 independent runs with varying random seed values are performed to obtain the mean testing $L^{2}$ error.
Figure 6: Generalization performance comparison on 3D HIT data at a resolution of $64^{3}$ and rank $r=5$, with a 2D slice of $u_{y}$. Top: $u_{y}$. Bottom: absolute error of $u_{y}$ between the reconstructed field and ground truth.
4.2 Koopman Forecasting Performance
We compare the temporal extrapolation performance of the different methods in conjunction with the Koopman forecast model. It can be seen from fig. 7, that our approach achieves an order of magnitude lower testing error in comparison to the other methods for the traveling wave case, with simple hybrid performing similarly to POD as seen in some of the previous examples. For cylinder flow, both the simple hybrid and the proposed framework produce an order of magnitude lower testing error compared to POD and AE, and our framework further reduces the error by a factor of 2 compared to the simple hybrid. To better understand the prediction from these frameworks, fig. 8 shows the snapshot of last 1000 timesteps of the traveling wave for the training and testing regime and the prediction from the different methods. Prediction from POD and Simple Hybrid shows stationary dynamics. AE predicts a superposition of two different waves. The prediction of the proposed approach matches very well with the ground-truth data. The POD and simple hybrid approaches appear to be constrained by the expressivity of the model, preventing them from identifying a low-dimensional embedding of $r=2$ that effectively represents the system. This is a known limitation of POD for such problems. In this case, the simple hybrid was observed to be dominated by the POD component, resulting in an expressivity similar to that of POD. Although AE outperforms POD and Simple hybrid , its performance was found to be constrained by the optimization process, often getting stuck in some local minima. Meanwhile, the proposed framework effectively captures the mapping to and from the latent space and mitigates the optimization challenges, since it is initialized with POD values at the start of training. The performance of the frameworks on the training data indicates that our proposed framework effectively learns both the spatial and temporal characteristics of the dataset. This capability extends to the testing regime, demonstrating that our model generalizes well. For the cylinder flow dataset fig. 9, the proposed framework demonstrates superior generalization capabilities compared to alternative methods. Although the AE and simple hybrid approaches achieve performance comparable to our framework in the training regime, they exhibit notable degradation when the dynamics are forecast beyond the training horizon. In contrast, our model maintains high-fidelity reconstructions. This further emphasizes the significant improvement that can be made through the introduction of a few useful weighting parameters.
Figure 7: Generalization performance of the Koopman decoder model on traveling wave and flow over cylinder dataset. Latent space rank of 2 and 4 was used for the traveling wave cylinder case respectively. Different methods for dimensionality reduction are represented by different colors. Y axis represents the Mean Testing $L^{2}$ error.
Figure 8: Comparison of solution state among the different method for traveling wave for a latent space rank of 2. The training data consists of the first 50,000 timesteps and the data for the remaining 50000 timesteps are held out for testing. Left: Last 1000 snapshots within the training regime. Right: Last 1000 snapshots within the testing regime.
Figure 9: Comparison of solution state among the different method for flow over cylinder for a latent space rank of 4. The training data consists of the first 50 timesteps and the data for the remaining 100 timesteps are held out for testing. Left: Final snapshot within the training regime (timestep 50). Right: Final snapshot within the testing regime (timestep 150).
4.3 Surrogate Modeling Performance
We evaluate POD, AE, Simple Hybrid, and the proposed Learnable Weighted Hybrid for time-dependent PDE surrogate modeling, where dimensionality reduction is combined with LSTM-based time series prediction to assess the impact of reduced latent representations on system dynamics forecasting. The mean test error $L^{2}$ for the different methods in the three PDEs considered for surrogate modeling is shown in fig. 10. The hatched portion denotes the error contribution from the time series modeling, and the solid region represents the error due to dimensionality reduction. In all the cases, it can be seen that our proposed approach maintains a superior performance in comparison to the other methods. We obtain almost an order-of-magnitude reduction in generalization error for the 1D VB. For 2D Shallow Water and 3D VB, the simple hybrid approach does not show much improvement over POD, while AE has the highest testing error.
Figure 10: Generalization performance of the surrogate models on the three PDEs: 1D Viscous Burgers’, 2D Shallow Water and 3D Viscous Burgers’. We use a latent space rank of 2 for the 1D Viscous Burgers’ case and 6 for the 2D Shallow Water and 3D Viscous Burgers’ case. Different methods for dimensionality reduction are represented by different colors. Y axis represents the Mean Testing $L^{2}$ error. The hatched portion denotes the error contribution from the LSTM model and solid region denotes the error contribution purely from dimensionality reduction.
It is to be noted that error from dimensionality reduction dominates over the error introduced by system dynamics modeling by orders of magnitude. This highlights the critical importance of constructing high quality reduced representations, as any downstream application in reduced order modeling like forecasting, control, or optimization is inherently limited by the accuracy of the latent space.
To further illustrate the effectiveness of the proposed approach, we compare the reconstructed solutions for the 1D VB problem at the final time of 2s for a $Re$ value of 2450 in fig. 11. In such high Reynolds number scenarios, the solution exhibits the formation and propagation of shocks over time, making accurate reduced-order modeling particularly challenging. It can be noted that methods like AE and Simple Hybrid exhibit high-frequency oscillations in space, particularly near the shock, whereas our approach maintains a smooth profile that closely conforms to the ground truth. POD solution although smooth but doesn’t capture the location and profile of the shock accurately. This demonstrates that our method is well-suited for handling convection dominated scenarios.
Figure 11: Generalization performance comparison on 1D Viscous Burgers’ data for a latent space rank of 2. The figure depicts the solution state at the final time for a trajectory with $Re=2450$. A zoomed in view near the shock location is provided for better visualization.
Next, we analyze the solution at the final time of 0.5s for the 2D Shallow Water case in fig. 12. Our approach effectively captures both large-scale and small-scale flow features within the solution domain. In contrast, POD and the Simple Hybrid approach primarily recover only the large-scale structures, while AE reconstruction contains non-sharp features. Similar comparison for the 3D VB can be found in fig. 13. While the predictions from our framework do not perfectly match the ground truth, they exhibit superior physical consistency and align more closely with the true solution than other methods, despite using only six latent dimensions. The other methods tend to predict an overly positive value for the velocities and generates totally nonphysical flow features.
Figure 12: Generalization performance comparison on 2D Shallow Water data for a latent space rank of 6. The figure depicts the solution state at the final time for a testing trajectory. Each column corresponds to a particular state variable being compared.
Figure 13: Generalization performance comparison on 3D VB data for a latent space rank of 6. The figure depicts the solution state at the final time for a testing trajectory. Each column corresponds to a particular state variable being compared.
4.4 Noise Robustness
Recent studies [28] in the deep learning community show that the sharpness of the minima, which describes the sensitivity of model loss with respect to perturbations in the model parameters, is a promising quantity that correlates with the generalization performance of deep networks. Let $\mathcal{D}_{\textrm{train}}=\left\{(s_{1},g_{1}),\ldots,(s_{n},g_{n})\right\}$ be the training data, that is, the set of features $s$ and target $g$ pairs, and $\ell_{i}(\mathbf{\Theta})$ be the loss of an NN model parametrized by weights $\mathbf{\Theta}$ and evaluated at the $i$th training sample point $\left(s_{i},g_{i}\right)$. Afterwards, the sharpness on a set of points $\mathcal{D}\subseteq\mathcal{D}_{\text{train}}$ can be defined as [29]: $\mathcal{S}(\mathbf{\Theta};\mathcal{D})\triangleq\max_{\|\delta\|_{2}\leq\rho}\frac{1}{|\mathcal{D}|}\sum_{i:\left(s_{i},g_{i}\right)\in\mathcal{D}}(\ell_{i}(\mathbf{\Theta}+\delta)-\ell_{i}(\mathbf{\Theta}))$ where $\delta$ is the perturbation introduced on the trainable parameters $\mathbf{\Theta}$ and $\rho$ refers to the perturbation radius. For example, table 4 shows the sharpness and reconstruction error for a particular training instance of 3D HIT data with grid points $32^{3}$ and rank 5. Our proposed framework has 1000 times less sharpness compared to the AE and simple hybrid approach. Since the sharpness of the minima is related to the resilience of the model in the presence of noisy data, we evaluate the reconstruction performance of AE, a learnable weighted simple hybrid framework under noisy testing data. Models trained on noise-free data are tested on data that contain random normal noise with zero mean and standard deviation equal to $10\%$, $20\%$ and $30\%$ of the maximum velocity magnitude superimposed over the flow field.
Method Sharpness Reconstruction $L^{2}$ Error Under Varying Noise Level (of Maximum Velocity)
No Noise 10% 20% 30%
AE $16.69\pm 0.9$ 9.139.13{9.13}9.13 11.50 25.95 41.20
Simple Hybrid $22.61\pm 0.82$ 18.7618.76{18.76}18.76 22.26 36.93 58.12
Learnable Weighted Hybrid (Ours) $\mathbf{0.015}\pm\mathbf{0.001}$ $\mathbf{7.17}$ $\mathbf{8.12}$ $\mathbf{9.33}$ $\mathbf{9.43}$
Table 4: Sharpness and reconstruction $L^{2}$ error for three deep autoencoders with rank $r=5$ on the 3D HIT dataset at a resolution of $32^{3}$ under varying noise levels in the test input.
Figure 14: Comparison of the reconstruction performance of three different deep autoencoder frameworks on 3D HIT dataset at resolution $32^{3}$ and rank $r=5$. Up: noise free. Down: 30% noise level.
The reconstruction error for all models is shown in the same table. As the noise level increases, there is a drastic increase in the testing error for both the AE and the simple hybrid method, while our approach stays the same, demonstrating the robustness of this framework against noise in the unseen test data. A 2D slice of $u_{y}$ with and without 30% noise level is further visualized in fig. 14, which shows the improved robustness of our approach over other methods.
5 Conclusions
In this work, we present a novel deep autoencoder framework that demonstrates convergence properties akin to SVD. By incorporating a learnable weighted average between SVD and vanilla deep autoencoders (either feedforward or convolutional), our approach achieves SVD-like convergence as the rank increases. We validate the effectiveness of this framework on pure reconstruction tasks using two challenging chaotic PDE datasets: the 1D Kuramoto-Sivashinsky and the 3D homogeneous isotropic turbulence. The results show that our learnable weighted hybrid autoencoder consistently achieves the lowest testing error and exhibits superior robustness to noisy data compared to other methods such as POD, vanilla deep autoencoders, and simple hybrid autoencoders. Remarkably, we find that our proposed approach leads to a minimum with a sharpness that is a thousand times smaller than that of other deep autoencoder frameworks. In addition, we demonstrate that utilizing our proposed framework in tandem with time series prediction models, we can achieve superior performance for surrogate modeling of time-dependent PDE over other approaches. Our framework is also capable of capturing dynamics of the system with strong discontinuities without spurious oscillations in the solution. This highlights its potential for robust and generalizable representation learning in complex PDE systems.
CRediT authorship contribution statement
Nithin Somasekharan: Data Curation (lead), Formal Analysis (lead), Investigation (lead), Software (lead), Visualization (lead), Writing – Original Draft Preparation (lead). Shaowu Pan: Conceptualization (lead), Funding Acquisition (lead), Methodology (lead), Supervision (lead), Writing – Review & Editing (lead), Project Administration (lead).
Acknowledgements
This work was supported by U.S. Department of Energy under Advancements in Artificial Intelligence for Science with award ID DE-SC0025425. The authors thank the Center for Computational Innovations (CCI) at Rensselaer Polytechnic Institute (RPI) for providing computational resources during the early stages of this research. Numerical experiments are performed using computational resources granted by NSF-ACCESS for the project PHY240112 and that of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility using the NERSC award NERSC DDR-ERCAP0030714.
Appendix A Detailed comparison of numerical results
We summarize the detailed results of all numerical experiments conducted in this work in tables 5 and 6. Upon publication, the code will be available at https://github.com/csml-rpi/deep-ae-with-svd-convergence.
Grid Rank Mean Train Error, $\mu$(±$\pm\sigma$) Mean Test Error, $\mu$(±$\pm\sigma$) Total Number of Model Parameters
POD AE Simple Learnable Weighted POD AE Simple Learnable Weighted POD AE Simple Learnable Weighted
Hybrid Hybrid (ours) Hybrid Hybrid (ours) Hybrid Hybrid (ours)
512 50 2.84e-01 2.54e-03 2.41e-04 1.62e-04 3.06e-01 2.73e-03 2.55e-04 1.92e-04 25,600 113,162 138,762 138,813
($\pm$ 1.92e-03) ($\pm$1.42e-03) ($\pm$6.07e-05) ($\pm$9.47e-06) ($\pm$4.80e-03) ($\pm$1.49e-03) ($\pm$5.96e-05) ($\pm$1.03e-05)
60 6.89e-02 1.68e-03 1.88e-04 1.59e-05 7.50e-02 1.76e-03 1.94e-04 1.96e-05 30,720 138,092 168,812 168,873
($\pm$ 4.92e-04) ($\pm$5.84e-04) ($\pm$7.35e-05) ($\pm$7.32e-07) ($\pm$1.47e-03) ($\pm$5.75e-04) ($\pm$7.38e-05) ($\pm$2.94e-06)
70 1.83e-02 1.75e-03 1.93e-04 2.73e-06 2.03e-02 1.79e-03 1.97e-04 3.55e-06 35,840 163,822 199,662 199,733
($\pm$ 1.37e-04) ($\pm$1.14e-03) ($\pm$1.26e-04) ($\pm$3.95e-07) ($\pm$3.67e-04) ($\pm$1.10e-03) ($\pm$1.28e-04) ($\pm$9.03e-07)
80 4.59e-03 2.07e-03 1.55e-04 1.02e-06 5.16e-03 2.22e-03 1.58e-04 1.39e-06 40,960 190,352 231,312 231,393
($\pm$ 2.95e-05) ($\pm$1.68e-03) ($\pm$4.62e-05) ($\pm$1.05e-07) ($\pm$8.78e-05) ($\pm$1.75e-03) ($\pm$4.72e-05) ($\pm$5.40e-07)
1024 50 3.18e-01 6.48e-03 3.01e-04 2.53e-04 3.41e-01 6.95e-03 3.20e-04 2.98e-04 51,200 216,074 267,274 267,325
($\pm$ 2.28e-03) ($\pm$2.53e-03) ($\pm$6.88e-05) ($\pm$2.29e-05) ($\pm$5.12e-03) ($\pm$2.49e-03) ($\pm$6.89e-05) ($\pm$2.29e-05)
60 8.75e-02 6.10e-03 2.14e-04 2.89e-05 9.53e-02 6.19e-03 2.22e-04 3.61e-05 61,440 261,484 322,924 322,985
($\pm$ 4.46e-04) ($\pm$7.01e-03) ($\pm$6.56e-05) ($\pm$3.29e-06) ($\pm$1.11e-03) ($\pm$6.43e-03) ($\pm$6.54e-05) ($\pm$5.84e-06)
70 2.21e-02 3.76e-03 2.25e-04 5.33e-06 2.46e-02 4.08e-03 2.30e-04 7.10e-06 71,680 307,694 379,374 379,445
($\pm$ 9.31e-05) ($\pm$2.64e-03) ($\pm$1.31e-04) ($\pm$1.31e-06) ($\pm$3.32e-04) ($\pm$2.72e-03) ($\pm$1.30e-04) ($\pm$2.38e-06)
80 5.86e-03 3.69e-03 2.23e-04 1.94e-06 6.58e-03 4.00e-03 2.26e-04 2.68e-06 81,920 354,704 436,624 436,705
($\pm$ 3.16e-05) ($\pm$2.34e-03) ($\pm$1.07e-04) ($\pm$1.54e-07) ($\pm$9.69e-05) ($\pm$2.34e-03) ($\pm$1.06e-04) ($\pm$8.49e-07)
2048 50 2.92e-01 1.62e-02 2.74e-04 2.36e-04 3.15e-01 1.82e-02 2.98e-04 2.80e-04 102,400 421,898 524,298 524,349
($\pm$ 2.34e-03) ($\pm$1.81e-02) ($\pm$7.71e-05) ($\pm$2.03e-05) ($\pm$5.32e-03) ($\pm$1.88e-02) ($\pm$8.97e-05) ($\pm$3.50e-05)
60 7.20e-02 9.93e-03 1.65e-04 3.39e-05 7.84e-02 1.14e-02 1.76e-04 4.35e-05 122,880 508,268 631,148 631,209
($\pm$ 4.95e-04) ($\pm$4.29e-03) ($\pm$2.48e-05) ($\pm$2.63e-06) ($\pm$1.31e-03) ($\pm$4.79e-03) ($\pm$2.31e-05) ($\pm$1.05e-05)
70 1.80e-02 7.65e-03 2.52e-04 1.21e-05 1.99e-02 9.18e-03 2.59e-04 1.61e-05 143,360 595,438 738,798 738,869
($\pm$ 9.85e-05) ($\pm$3.15e-03) ($\pm$7.35e-05) ($\pm$6.28e-07) ($\pm$3.20e-04) ($\pm$3.47e-03) ($\pm$7.46e-05) ($\pm$4.49e-06)
80 4.72e-03 9.62e-03 3.07e-04 9.23e-06 5.34e-03 1.12e-02 3.13e-04 1.23e-05 163,840 683,408 847,248 847,329
($\pm$ 3.99e-05) ($\pm$6.61e-03) ($\pm$2.34e-04) ($\pm$3.11e-07) ($\pm$1.54e-04) ($\pm$6.77e-03) ($\pm$2.30e-04) ($\pm$2.95e-06)
Table 5: Summary of training $L^{2}$ error, testing $L^{2}$ error, and the number of parameters for the four models trained on data with varying grid resolutions and ranks $r$ pertaining to 1D K-S case. The standard deviation of the error is indicated in parenthesis. The number of parameters for the hybrid approaches includes non-trainable POD parameters, which remain fixed throughout the optimization process and are not trainable.
Grid Rank Mean Train Error, $\mu$(±$\pm\sigma$) Mean Test Error, $\mu$(±$\pm\sigma$) Total Number of Model Parameters
POD AE Simple Learnable Weighted POD AE Simple Learnable Weighted POD AE Simple Learnable Weighted
Hybrid Hybrid (ours) Hybrid Hybrid (ours) Hybrid Hybrid (ours)
$16^{3}$ 5 2.01e+01 1.05e+00 1.35e+00 1.76e+00 2.04e+01 5.74e+00 5.98e+00 5.48e+00 61,440 688,348,168 688,409,608 688,409,616
($\pm$ 0.00) ($\pm$1.37e-01) ($\pm$1.13e-01) ($\pm$9.31e-02) ($\pm$0.00) ($\pm$3.13e-01) ($\pm$1.09e-01) ($\pm$6.44e-02)
10 1.41e+01 9.71e-01 6.85e-01 9.02e-01 1.47e+01 5.53e+00 5.17e+00 5.22e+00 122,880 688,368,653 688,491,533 688,491,546
($\pm$ 0.0) ($\pm$2.75e-01) ($\pm$6.05e-02) ($\pm$4.04e-02) ($\pm$0.0) ($\pm$1.11e-01) ($\pm$3.20e-02) ($\pm$3.49e-02)
15 1.14e+01 8.76e-01 4.65e-01 6.35e-01 1.23e+01 5.47e+00 5.08e+00 5.20e+00 184,320 688,389,138 688,573,458 688,573,476
($\pm$ 0.0) ($\pm$4.36e-02) ($\pm$2.43e-02) ($\pm$3.55e-02) ($\pm$0.00) ($\pm$1.02e-01) ($\pm$3.22e-02) ($\pm$3.59e-02)
20 9.71e+00 8.96e-01 3.07e-01 5.06e-01 1.09e+01 5.49e+00 4.85e+00 5.21e+00 245,760 688,409,623 688,655,383 688,655,406
($\pm$ 0.00) ($\pm$5.07e-02) ($\pm$1.90e-02) ($\pm$2.83e-02) ($\pm$0.00) ($\pm$6.92e-02) ($\pm$3.01e-02) ($\pm$3.14e-02)
$32^{3}$ 5 2.02e+01 9.15e+00 1.84e+01 5.56e+00 2.04e+01 1.05e+01 1.87e+01 7.16e+00 491,520 688,505,864 688,997,384 688,997,392
($\pm$ 0.00) ($\pm$1.10e+01) ($\pm$2.61e-01) ($\pm$1.03e-01) ($\pm$0.00) ($\pm$1.05e+01) ($\pm$2.34e-01) ($\pm$7.28e-02)
10 1.42e+01 1.52e+01 1.23e+01 3.98e+00 1.47e+01 1.63e+01 1.32e+01 6.50e+00 983,040 688,669,709 689,652,749 689,652,762
($\pm$ 0.00) ($\pm$1.74e+01) ($\pm$1.45e-01) ($\pm$6.67e-02) ($\pm$0.00) ($\pm$1.66e+01) ($\pm$1.27e-01) ($\pm$5.38e-02)
15 1.15e+01 1.46e+01 9.18e+00 2.90e+00 1.22e+01 1.59e+01 1.06e+01 6.18e+00 1,474,560 688,833,554 690,308,114 690,308,132
($\pm$ 0.00) ($\pm$1.76e+01) ($\pm$1.81e-01) ($\pm$5.54e-02) ($\pm$0.00) ($\pm$1.68e+01) ($\pm$1.55e-01) ($\pm$4.65e-02)
20 9.78e+00 1.10e+01 6.91e+00 2.21e+00 1.08e+01 1.26e+01 8.86e+00 5.97e+00 1,966,080 688,997,399 690,963,479 690,963,502
($\pm$ 0.00) ($\pm$1.52e+01) ($\pm$2.84e-01) ($\pm$4.70e-02) ($\pm$0.00) ($\pm$1.44e+01) ($\pm$1.55e-01) ($\pm$3.96e-02)
$64^{3}$ 5 2.02e+01 2.89e+01 1.95e+01 9.00e+00 2.04e+01 2.89e+01 1.96e+01 9.53e+00 3,932,160 689,767,432 693,699,592 693,699,600
($\pm$ 0.00) ($\pm$1.94e+01) ($\pm$1.53e-01) ($\pm$9.62e-02) ($\pm$0.00) ($\pm$1.91e+01) ($\pm$1.48e-01) ($\pm$9.27e-02)
10 1.42e+01 2.50e+01 1.35e+01 7.01e+00 1.46e+01 2.52e+01 1.40e+01 7.95e+00 7,864,320 691,078,157 698,942,477 698,942,490
($\pm$ 0.00) ($\pm$2.00e+01) ($\pm$1.22e-01) ($\pm$7.07e-02) ($\pm$0.00) ($\pm$1.96e+01) ($\pm$1.18e-01) ($\pm$6.18e-02)
15 1.15e+01 1.65e+01 1.08e+01 5.59e+00 1.22e+01 1.69e+01 1.16e+01 7.01e+00 11,796,480 692,388,882 704,185,362 704,185,380
($\pm$ 0.00) ($\pm$1.65e+01) ($\pm$1.79e-01) ($\pm$3.82e-02) ($\pm$0.00) ($\pm$1.62e+01) ($\pm$1.69e-01) ($\pm$3.63e-02)
20 9.78e+00 1.94e+01 9.13e+00 4.54e+00 1.08e+01 1.97e+01 1.02e+01 6.45e+00 15,728,640 693,699,607 709,428,247 709,428,270
($\pm$ 0.00) ($\pm$1.82e+01) ($\pm$9.70e-02) ($\pm$3.73e-02) ($\pm$0.00) ($\pm$1.78e+01) ($\pm$9.29e-02) ($\pm$3.39e-02)
Table 6: Summary of training $L^{2}$ error, testing $L^{2}$ error error, and number of parameters for the four models trained on data with varying resolutions $N$ and ranks $r$ pertaining to 3D HIT case.